# Towards AI Accountability Infrastructure: Gaps and Opportunities in AI Audit Tooling

### v1

@@ 3.2. Tool Taxonomy
Initial search. To taxonomize the landscape of available audit tools, we first developed an initial list of tools and tool-building organizations by searching for tools mentioned in academic audit studies, news articles, government reports and frameworks, white papers from civil society organizations, law firm reports, case files, and existing lists of tools such as (hickock_ethical_nodate, 46) (see Appendix A.1 for details). Our initial search was conducted in August 2022. We also included specific tools and sources that were mentioned in our interviews with practitioners. After we had collected an initial list of 148 tools, we developed an initial taxonomy by clustering tools into 21 initial categories based on their intended or actual uses in AI audit work.
@@
Theoretical sampling. Next, we expanded our initial set of tools with two kinds of additional theoretical sampling. With targeted keyword searches on English Google and GitHub, we searched explicitly for additional tools in areas where we had fewer examples until theoretically fresh examples of tools ceased to arise. [^3] In this stage, we aimed for theoretical saturation, in the style of grounded theory (charmaz_constructing_2014, 21). Our search queries were the category descriptors—either from our initial categories or from descriptors used by tools already collected—alone or combined with terms like “audit tool” or “responsible AI” (see Appendix A.2). We also expanded our list of sources based on our initial taxonomy (e.g., for our initial “participatory” category, searching for tools mentioned in the Participatory Approaches to Machine Learning Workshop). We also followed links and references in our initial sample of tools to identify additional, similar tools (snowball sampling). Between August and October 2022, we added 181 tools with these methods, and we continued to update the database with 61 more examples through September 2023.
@@
These searches surfaced new examples that forced us to expand and re-define our categories or create new ones. We iteratively revised our taxonomy twice more to accommodate new examples and to clarify or expand our categories, grouping clusters as more general themes appeared. Our final taxonomy groups tools into 30 main categories with 27 subcategories, sorted into a final set of 7 “stages” of the tool-assisted audit process (Fig. 1). The resulting taxonomy [^4] Readers can view and contribute to a full interactive version of our taxonomy and dataset at tools.auditing-ai.com. is grounded in the properties of the tools we found and shaped by our interviews with and our own experiences as AI audit practitioners.
@@
Table 2. High-level description of the tool taxonomy categories (visit tools.auditing-ai.com for an interactive visualization).
@@
Education / Awareness (community education, visioning), Incident Reporting (incident databases, intake forms, bug bounties, hotlines), Target Identification (algorithm visibility)
45
Help auditors identify and prioritize audit targets and harms to investigate.
ACLU Wa.’s Algorithm Equity Toolkit, AI Incident Database, Algorithm Tips
Standards Id. & Mgmt.
@@
Goal Articulation (principle statements, standards formulation), Self-Assessment (checklists, grading), Documentation (single stage, continuous, licenses), Regulatory Awareness (discovery, monitoring), Methods Design, Participatory Standards-Setting
194
Help auditors identify and formulate principles and norms to guide their investigations.
AI-RFX Procurement Framework, Microsoft’s AI Fairness Checklist, Model Cards (mitchell_model_2019, 68), Queensland’s Community Engagement Toolkit (queensland_government_community_2017, 77), Community Jury
Transparency Infrastructure
Structured/API Access, Secure & Private Sharing (federated learning), Model/Data Exchange
13
Help auditors interact with and analyze proprietary information about the data or model with centralized infrastructure.
@@
NIST’s Face Recognition Vendor Test (ngan_ongoing_2020, 71), Google AI Test Kitchen (warkentin_join_2022, 107), Airbnb’s Project Lighthouse (airbnb_new_2020, 3)
Data Collection
Field Data Collection (scraping, donation, interviews/surveys, compelled disclosure), Bot Deployment, Simulation
43
Help auditors collect information about a model’s interactions with its subjects.
Mozilla’s YouTube Regrets (mozilla_foundation_youtube_2021, 69), Tracking Exposed, Selenium, Meta’s Web-Enabled Simulation (ahlgren_wes_2020, 2)
Performance Analysis
Accuracy Evaluation (A/B testing, benchmarks, adversarial testing, monitoring), Explainability (models, training data), Fairness, Qualitative Analysis
126
Help auditors evaluate and explain model behavior through the calculation of performance metrics.
@@
Weights & Biases, Meta’s DynaBench, Foolbox, Fairlearn, IBM’s AI Fairness 360, Hugging Face’s ROOTS (piktus_roots_2023, 75), Google PAIR’s Language Interpretability Tool
Advocacy
Resistance, Community Spaces, Legal Search, Organizing
14
Help organize community action and other accountability measures in response to discovered harms.
Gigbox, Para, Adnauseam, Benefits Tech Advocacy Hub
Audit Communication
Dataset Visualization
2
@@
Help auditors communicate the results of an audit to a broader audience.
Google PAIR’s FACETS
3.3. Landscape analysis
To analyze the qualities of tools across our taxonomy,555 All the code for our analysis and resulting plots—as well as instructions for obtaining supplemental data—can be accessed at github.com/ryansteed/oat-analysis. we also manually labeled each tool with several tags describing the tool’s documentation and function: license (open-source or proprietary); organization type (for-profit, non-profit, government, or academic); intended audit target (automated decision system, online platforms, large pre-trained online platforms autonomous vehicles, and/or other); intended user (internal and/or external); and format (e.g. API, software product, code/data repository, white paper, and/or other). One author created the labels and at least one other author reviewed each label for agreement.
We also supplemented our dataset with data from Crunchbase (noauthor_crunchbase_nodate, 25), a platform for tracking funding, employment, revenue and other data for technology ventures, and Github (noauthor_github_nodate, 42), a platform for hosting and developing software. From Github, we scraped repository activity—primarily the number of forks, stars, and issues—for the 98 tools with Github repositories in our dataset. 187 of the 312 organizations in our dataset (accounting for 247 tools) have Crunchbase entries, which we manually identified using fuzzy matching and the Crunchbase search feature; 171 include estimated employee counts. Of the 129 entries that are not for universities or government agencies, 88 include revenue estimates. We also collected total venture funding (adjusted to U.S. dollars) for 47 firms out of the 110 firms that are still private (i.e., have not undertaken an initial public offering). Additionally, we used the Google Scholar API to annotate each academic reference of an identified tool with the most recent available citation count.
We analyzed the distribution of these supplemental variables across our taxonomy and include our findings in §4. Detailed results, including all figures summarizing all our quantiative analysis, can be found in Appendix C.
3.4. Reflections
We consider our own cultural and professional perspectives throughout the interviews and our analysis (charmaz_constructing_2014, 21). In particular, we view our position as both external to prominent AI developers and deployments, but still situated within the Western AI industry—our project was financially supported by a prominent U.S.-based foundation, all the authors are either graduate students or graduates of well-funded universities in the U.S. and Europe. Three have some experience studying or developing AI systems at U.S. tech companies. Because we drew our initial sources from our own fieldwork as audit practitioners and used English search engines for theoretical sampling, our dataset consists primarily of English language tools from from Western organizations, and our taxonomy reflects our own particular position [^6] We also attempted to run translated queries in non-English search engines (e.g., Baidu) and to add regional keywords (e.g., African) to our searches, but we were unable to find any additional tools with initial tests of these methods. It thus seems likely that non-English, non-Western tools exist that our theoretical sampling was unable to identify..
While analyzing our data and presenting the results, we considered how our positions might impose certain framings on the data or close off possible lines of inquiry. For example, we recognize that our analysis is primarily scoped to the U.S. and the E.U. and may not be representative of the global AI auditing landscape or appropriate for or informed by other contexts. We wrote this paper with the goal of promoting holistic accountability for algorithm-enabled harms. In particular, we struggled to avoid a techno-solutionist framing in our conclusions, given that our project focuses on toolkits that often fail to address non-technical dimensions of audit work (wong_seeing_2023, 111). We intentionally defined tools as resources more broadly and attempted to leave space for non-technical solutions in our analysis and discussion, and we hope that future work will expand and look outside of the mostly technical, Western solutions analyzed in our work.
4. Results
Our survey of AI audit tools revealed a wide-ranging landscape of resources built by a variety of academic, for-profit, non-profit, and government organizations for a variety of purposes. A full interactive version of our dataset can be found at tools.auditing-ai.com, and further details about our analysis can be found at github.com/ryansteed/oat-analysis.
We identified seven main stages of tool-assisted auditing, depicted in Figure 1 and summarized in Table 2. Our survey surfaced many tools similar to those traditionally studied and developed in academic work—such as tools for Performance Analysis (particularly fairness evaluation) and Standards Identification & Management (such as principle statements or checklists). However, we also surfaced tools designed to support with other less-studied elements of AI auditing—such as Harms Discovery, Data Collection, and Transparency Infrastructure. For example, a few tools are uniquely focused on ensuring that audits result in accountability through facilitating Audit Communication and Advocacy. In part because our survey included only publicly listed tools, many of these tools (77.9%) were freely available and open-source, though we do include many proprietary or internal tools that are publicly advertised or publicly listed.
Our interviews with practitioners told a more complicated story. While many tools exist to aid in AI auditing, practitioners found the current array of resources inadequate in many ways. In particular, though many tools existed to aid in the evaluation of AI systems, current tooling did not always contribute to the ultimate aim of accountability. Despite the numerous tools available, practitioners still struggled to involve affected stakeholders, apply consistent and holistic standards and methods, access high-quality data about system behavior, ensure audit integrity, and collaborate across disciplines. And despite theoretical sampling, our survey of AI audit tools ($N=390$) revealed mostly tools for evaluation—particularly tools for Standards Identification & Management ($N=184$) and Performance Analysis ($N=116$). Tools for stages of the audit process crucial to accountability—Audit Communication ($N=2$), Harms Discovery ($N=43$), Advocacy ($N=13$), and model/data transparency ($N=12$)—are much less common. Figure 2 depicts the number of tools in each stage and category of our taxonomy.
@@ 4. Results
In this section, we detail the main tool-related challenges practitioners faced related to each stage in tool-assisted AI auditing, and discuss the implications of our findings for the practice and study of AI auditing. Note that we did not explicitly ask interview participants about the tool-assisted AI audit stages, but after interviews, categorize their interview responses within the format of our existing AI audit tool taxonomy, arranged by the seven stages (Fig. 1).
Figure 2. Number of tools in each category of our taxonomy, grouped by type of organization.
@@
Figure 2. Number of tools in each category of our taxonomy, grouped by type of organization.
@@ 4.1. Harms Discovery
A central task for AI auditing is identifying AI systems that should be subject to scrutiny and identifying their potential harms. This task can be especially difficult for external auditors who may not know where AI systems are in use or what their impacts might be. Tools for Harms Discovery ($N=43$) help identify and select targets for audits and support the identification, characterization, and prioritization of potential harms to investigate. This category includes tools for Education & Awareness (to engage affected stakeholders in articulating harms), Incident Reporting (to gather reports of algorithmic harms from users and the public, e.g., through bug bounties), and Target Identification (e.g., the Algorithm Tips database contains a list of deployed systems in the U.S. that could be investigated).
Compared to other stages of our taxonomy (Fig. 2), nonprofits contributed especially to creating and maintaining tools for Harms Discovery (e.g., databases for documenting AI incidents (charlie_pownall_ai_2021, 20)) and Data Collection (e.g., scraping or data donation (devos_toward_2022, 32)).
@@
Facilitating more participatory audits. Auditors recognized that to comprehensively identify AI-related harm, they must engage with those directly impacted by the AI deployments—though they did not necessarily have the tools to do so. Many participants recognized the importance of ongoing user engagement, collaborative feedback, and a focus on community representation in performing thorough and representative audits. Participation in harms discovery had two main benefits for auditors. First, participation helps auditors anticipate a broader range of possible harms: “Different types of biases are going to manifest, and accordingly it requires … diverse groups from society, to understand their experiences and expectations in these settings and how they can be impacted” (P20). Second, participation helped make auditing more “context-dependent” and inclusive by providing thorough understanding of how an AI system intersects with impacted groups. Participation could also help instill confidence in the audit process among these groups and foster engagement with subsequent accountability efforts. Our tool survey did surface quite a few tools for participatory incident reporting—such as the bug bounty platform HackerOne, which Twitter used to crowd audit its image cropping algorithm (twitter_twitter_2021, 102), or the AI Incident Database, a collection of reports of AI harms. However, fewer tools were designed specifically for identifying and collaborating with affected users, such as the American Civil Liberties Union (ACLU) of Washington’s Algorithmic Equity Toolkit (barghouti_algorithmic_2020, 8).
@@
Avoiding participation washing. However, participants also recognized the challenge of designing methodologies and tools without exploitation, tokenization, or other forms of “participation washing” (sloane_participation_2022, 93), highlighting the need for fair compensation and participatory algorithmic development, in addition to participatory auditing. One participant, for example, emphasized the importance of “an iterative process of doing longer-term, ongoing auditing or observation of algorithmic behavior, and then using that to feed into tweaks, or changes, or even big shifts in where and how [audit] outcomes are used” (P17).
@@
Implications. Limited access to information about AI systems poses a fundamental barrier to conducting comprehensive audits. Without a clear understanding of which AI systems require auditing, there is a risk of overlooking critical systems that have significant societal implications. While we did find multiple popular databases for recording and collating incidents of harm (such as the AI Incident Database), these databases mostly only record harms after they have already occurred, often rely on second-hand reports, and may not delve deeply into harms or impacts (turri_why_2023, 101).
@@
Several participants proposed mandating that corporations disclose AI systems, including information about model versions, anticipated use cases, expected users, and past audit results. Policymakers may also explore mechanisms for centralized, proactive documentation and mandatory, standardized incident reporting for both private firms and government agencies (turri_why_2023, 101)—and ensure that current federal AI transparency requirements are actually implemented (lawrence_bureaucratic_2023, 57). Additionally, enabling mechanisms such as Freedom of Information Act (FOIA) requests could facilitate access to information held by public institutions or government agencies regarding the use of AI systems in decision-making processes. Future work could also develop and study systems for fair, inclusive community participation in auditing, the path most often suggested by practitioners for identifying systems and their harms.

### v3

@@ 3.2. Tool Taxonomy
Initial search. To taxonomize the landscape of tools available to support AI audit work, we first developed an initial list of tools and tool-building organizations by searching for tools mentioned in a dataset of published audits from academic audit studies, news articles, government reports and frameworks, white papers from civil society organizations, law firm reports, and case files (birhane_ai_2024) as well as existing lists of tools such as (hickock_ethical_2023) (see Appendix C.1 for details). Our initial search was conducted in August 2022. We also included specific tools and sources that were mentioned in our interviews with practitioners. After we had collected an initial list of 143 tools, we developed an initial taxonomy by clustering tools into 21 initial categories based on their intended or actual uses in AI audit work. This approach served as a starting point for category development rather than as an exhaustive inventory.
@@
Theoretical sampling. Next, we expanded our initial set of tools with two kinds of additional theoretical sampling. With targeted keyword searches on English Google and GitHub, we searched explicitly for additional tools in areas where we had fewer examples until theoretically fresh examples of tools ceased to arise. [^7] In this stage, we aimed for theoretical saturation, in the style of grounded theory (charmaz_constructing_2014). Our search queries were descriptors from our initial categories or descriptors used by tools already collected—alone and combined with terms like “audit tool” or “responsible AI” (see Appendix C.2). We expanded our list of sources based on our initial taxonomy. (For example, to find tools in our initial “Participatory” category, we searched in the proceedings of participatory AI workshops). We also followed links and references in our initial sample of tools to identify additional, similar tools (snowball sampling). With these methods, we added 181 tools between August and October 2022, and we continued to update the dataset with 102 more tools through September 2024. In September 2024, we ran our search queries again for the top-level categories, adding 9 more tools.
@@
These searches surfaced new examples which we used to expand and re-define our categories. We iteratively revised our taxonomy twice more to accommodate new examples, integrate findings from the interview study, and clarify or expand our initial categories. We did not explicitly ask interview participants about these categories, but we did incorporate the tools they used and their descriptions of audit tool use while developing the taxonomy. Our final taxonomy groups tools into 30 main categories with 27 subcategories (Table 2) grounded in the properties of the tools we found and shaped by our interviews with and experiences as AI audit practitioners. We sorted these categories into 7 “stages” of the tool-assisted audit process (Fig. 1).
This stage of our search was designed to iteratively refine conceptual categories with the goal of generalizing from empirical descriptions to theoretical insights (lee_generalizing_2003). So while our search for tools was systematic, it was not exhaustive. The sample represented in our dataset does not include every tool that could support AI audits. Likewise, though our sample is designed to represent commonly used and commonly built tools in a theoretically representative set of categories, numeric descriptions of our sample may not reflect the statistical distribution of all AI audit tools used in practice. In particular, our theoretical sampling strategy deliberately over-represents some less common kinds of tooling (such as tools for advocacy). And because our search relied on public materials, we likely over-represent open tooling over proprietary tooling (though our dataset includes both). As a result, while our dataset reflects a structured and iterative search, it should not be viewed as a comprehensive or perfectly representative sample.
Landscape analysis. To analyze the qualities of tools across our taxonomy, we also manually labeled each tool with several tags describing the tool’s documentation and function, including license (open source or not), organization type (for-profit, non-profit, government, or academic), and other characteristics. One author created the labels and at least one other author reviewed each label for agreement. We also supplemented our dataset with funding & employment data from Crunchbase (accessed in September 2023), activity data from Github (September 2024), and citation data from Google Scholar API (September 2023). Detailed methods and additional analysis can be found in Appendix E, and a full interactive version of our dataset can be viewed at tools.auditing-ai.com.888 All the code for our analysis and resulting plots—as well as instructions for obtaining supplemental data—can be accessed at github.com/ryansteed/oat-analysis.
4. Results
While there exist many tools for aiding AI auditing, practitioners found existing tools inadequate in multiple ways. Practitioners struggled to independently access high-quality data about system behavior, apply consistent and holistic standards and methods, ensure audit integrity, involve affected stakeholders, and collaborate across disciplines.
In particular, though we found many tools for evaluating the performance of AI systems, current tooling did not always help practitioners reach their accountability goals. First, while our survey of AI audit tools ($N=435$), revealed a wide-ranging landscape of resources built by a variety of academic, for-profit, non-profit, and government organizations (Table 2), we primarily surfaced tools for evaluation, particularly tools for Standards Identification & Management ($N=206$) and Performance Analysis ($N=129$). Tools for other stages of the audit process crucial to accountability—Harms Discovery ($N=48$), Audit Communication ($N=8$), Advocacy ($N=14$), and model/data transparency ($N=12$)—were much less common in our sample.
@@ 4. Results
Table 2. High-level description of the tool taxonomy categories. (Visit tools.auditing-ai.com for an interactive visualization).
@@
Harms Discovery Education / Awareness (community education, visioning), Incident Reporting (incident databases, intake forms, bug bounties, hotlines), Target Identification (algorithm visibility) 48 Help auditors identify and prioritize audit targets and harms to investigate. ACLU Wa.’s Algorithm Equity Toolkit, AI Incident Database, Algorithm Tips
@@
Standards Identification & Mgmt. Goal Articulation (principle statements, standards formulation), Self-Assessment (checklists, grading), Documentation (single stage, continuous, licenses), Regulatory Awareness (discovery, monitoring), Methods Design, Participatory Standards-Setting 206 Help auditors identify and formulate principles and norms to guide their investigations. AI-RFX Procurement Framework, Microsoft’s AI Fairness Checklist, Model Cards (mitchell_model_2019), Queensland’s Community Engagement Toolkit (queensland_government_community_2017), Community Jury
@@
Transparency Infrastructure Structured/API Access, Secure & Private Sharing (federated learning), Model/Data Exchange 12 Help auditors interact with and analyze proprietary information about the data or model with centralized infrastructure. NIST’s Face Recognition Vendor Test (ngan_ongoing_2020), Google AI Test Kitchen (warkentin_join_2022), Airbnb’s Project Lighthouse (airbnb_new_2020)
Data Collection Field Data Collection (scraping, donation, interviews/surveys, compelled disclosure), Bot Deployment, Simulation 47 Help auditors collect information about a model’s interactions with its subjects. Mozilla’s YouTube Regrets (mozilla_foundation_youtube_2021), Tracking Exposed, Selenium, Meta’s Web-Enabled Simulation (ahlgren_wes_2020)
@@
Performance Analysis Accuracy Evaluation (A/B testing, benchmarks, adversarial testing, monitoring), Explainability (models, training data), Fairness, Qualitative Analysis 129 Help auditors evaluate and explain model behavior through the calculation of performance metrics. Weights & Biases, Meta’s DynaBench, Foolbox, Fairlearn, IBM’s AI Fairness 360, Hugging Face’s ROOTS (piktus_roots_2023), Google PAIR’s Language Interpretability Tool
@@
Audit Communication Dataset Visualization, Audit Reporting 8 Help auditors communicate the results of an audit to a broader audience. Google PAIR’s FACETS
Advocacy Organizing/Resistance, Community Spaces, Legal Search 14 Help organize community action and other accountability measures in response to discovered harms. Gigbox, Para, Adnauseam, Benefits Tech Advocacy Hub
Second, while we found many freely available and open source tools (77.9% of our dataset), auditors highlighted the messy, context-specific nature of their actual audit tool use: “Many approaches were not necessarily principled. They were quite ad hoc” (P20). Even when open source tools existed, auditors often preferred to build their own tooling solutions: “if we tried to use the existing stuff, it would just complicate that process” (P25). For some, existing tools were inadequate for the complexity and scale of the systems being evaluated: “Most often we try to use open source tools, but that’s very different than a data pipeline that… curates data on millions of [users] every day” (P3). In each stage of the audit process, auditors encountered practical challenges and development gaps between their needs and the landscape of available tools.
@@
In the remainder of this section, we detail the challenges practitioners faced in each stage, compare their experiences to the existing landscape of AI audit tools, and discuss the implications of our findings for the practice and study of AI auditing.
Figure 2. Number of tools in each category within each stage of our taxonomy, grouped by type of organization. Tools may be used in multiple stages. Note that the scales differ—the Standards and Performance Analysis stages contain many more tools than the others. Nonprofit and university/academic developers account for relatively more Harms Discovery and Data Collection tools. For-profit developers contribute relatively more Performance Analysis and Transparency Infrastructure tools.
[…]
@@ 4.1. Harms Discovery
Auditing an AI system first requires identifying the system that should be subject to scrutiny and identifying its potential harms. This task can be especially difficult for external auditors who may not know where AI systems are in use or what their impacts might be. Tools for Harms Discovery ($N=48$) help identify and select targets for audits and support the identification, characterization, and prioritization of potential harms to investigate. This category includes tools for Education & Awareness (to engage affected stakeholders in articulating harms), Incident Reporting (to gather reports of algorithmic harms from users and the public, e.g., through bug bounties or incident databases (charlie_pownall_ai_2021)), and Target Identification (e.g., the Algorithm Tips database contains a list of deployed systems in the U.S.). Compared to other stages of our taxonomy (Fig. 2), nonprofits contributed significantly to creating and maintaining these types of tools in our dataset (79.2% not-for-profit; see Fig. E.6).
@@
Facilitating more participatory audits. Auditors recognized that to comprehensively identify AI-related harms, they must engage with those directly impacted. Participation in harms discovery had two main benefits for auditors. First, participation helped auditors anticipate a broader range of possible harms: “Different types of biases are going to manifest, and accordingly it requires … diverse groups from society, to understand their experiences and expectations in these settings and how they can be impacted” (P20). Second, participation helped make audits more “context-dependent” and inclusive by providing thorough understanding of how an AI system interacts with impacted groups. Participation also helps instill confidence in the audit process and foster engagement with subsequent accountability efforts. Our tool survey surfaced some tools for participatory incident reporting. This includes the bug bounty platform HackerOne, which Twitter used to crowd audit its image cropping algorithm (twitter_twitter_2021), and the AI Incident Database, a collection of reports of AI harms. However, we found fewer tools designed specifically for identifying and collaborating with affected users, such as the American Civil Liberties Union (ACLU) of Washington’s Algorithmic Equity Toolkit (barghouti_algorithmic_2020).
@@
Avoiding participation washing. Participants also recognized the challenge of designing methodologies and tools without exploitation, tokenization, or other forms of “participation washing” (sloane_participation_2022), highlighting the need for fair compensation and participatory algorithmic development, in addition to participatory auditing. One participant emphasized the importance of “an iterative process of doing longer-term, ongoing auditing or observation of algorithmic behavior, and then using that to feed into tweaks, or changes, or even big shifts in where and how [audit] outcomes are used” (P17).
@@
Implications. Limited access to information about AI systems poses a fundamental barrier to conducting comprehensive audits. Without a clear understanding of which AI systems require auditing, there is a risk of overlooking critical systems that have significant societal implications. While we did find multiple popular databases for recording and collating incidents of harm (such as the AI Incident Database), these databases record harms after they have already occurred, often rely on second-hand reports, and may not delve deeply into causes of harms or impacts (turri_why_2023). [^9] Note: The creators of the AI and Algorithmic Incident and Controversies Repository dispute the characterization of their tool in turri_why_2023.
@@
Several participants proposed mandating that corporations disclose AI system use, including information about model versions, anticipated use cases, expected number of users, and past audit results. Researchers and policymakers may also explore mechanisms for centralized, proactive documentation and mandatory, standardized incident reporting for both private firms and government agencies (turri_why_2023). This ensures that current federal AI transparency requirements are actually implemented (lawrence_bureaucratic_2023). Additionally, leveraging mechanisms such as Freedom of Information Act (FOIA) requests could facilitate access to information held by public institutions or government agencies regarding the use of AI systems. Future work could also develop and study systems for fair, inclusive community participation in auditing, the path most often suggested by practitioners for identifying systems and their harms.